OpenStudy (anonymous):

Solve, using linear combination. 3x + y = 4 2x + y = 5 A. (–1, 7) B. (–3, 12) C. (0, 4) D. (–3, 13)

OpenStudy (anonymous):

a. (-1,7)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yvw :)

OpenStudy (anonymous):

i have a few more wold u help

OpenStudy (anonymous):

Use the graph to state the solution for the system. algebra unit 8 lesson 3 assessment question 2 x – y = –3 (line a) x + y = 3 (line b) A. (3, 0) B. (–3, 0) C. (0, 3) D. (0, –3)

OpenStudy (anonymous):

(3,0)

OpenStudy (anonymous):

hey @csscgirl do u know or wanna know how to change ur pic so its not a owl

OpenStudy (anonymous):

or the os owl

OpenStudy (anonymous):

sure :)

OpenStudy (texaschic101):

(0,3) is your answer

OpenStudy (anonymous):

ok so its easy just look at the to right corner and it says settings click it the in the about middle area there is ur pic and it says change picture

OpenStudy (anonymous):

thanks @texaschic101

OpenStudy (anonymous):

How many solutions are there to the following system of equations? 2x + 5y = –11 10x + 25y = –55 A. infinitely many B. 0 C. 2 D. 1

OpenStudy (anonymous):

oh thank you :)

OpenStudy (anonymous):

yep np B)

OpenStudy (anonymous):

a. infinitely many

OpenStudy (anonymous):

r u sure cause i dont wanna fail i got a f in math so im tryin to redo

OpenStudy (anonymous):

if you divide the whole thing by 5, you see that it is the same equation, so they are right on top of eachother.

OpenStudy (anonymous):

this is my last chance to pass

OpenStudy (texaschic101):

infinitely many is correct

OpenStudy (anonymous):

ooo ok

OpenStudy (anonymous):

thanks

OpenStudy (texaschic101):

anytime :)

OpenStudy (anonymous):

yep it is correct... use desmos.com/calculator to graph things, because it will help a lot, for this lesson or unit you are on, it will give you correct answers :)

OpenStudy (anonymous):

yvw

OpenStudy (anonymous):

How many solutions are there to the following system of equations? 3x + 7y = –10 15x + 35y = –45 A. 0 B. 1 C. infinitely many D. 2

OpenStudy (anonymous):

@csscgirl U STILL THERE

OpenStudy (texaschic101):

15x + 35y = -45 --- divide by 5 3x + 7y = - 9 and your other equations is 3x + 7y = -10 if your lines have the same slope but different y intercepts, then they are parallel lines that do not intersect. Therefore, there is no solution.

OpenStudy (anonymous):

so a thanks

OpenStudy (texaschic101):

sure thing :)

OpenStudy (anonymous):

Solve the following system of equations by linear combination: 2d + e = 8 d – e = 4 A. The solution is (5, –2). B. There is no solution. C. The solution is (4, 0). D. There are an infinite number of solutions.

OpenStudy (anonymous):

i thought it was c

OpenStudy (texaschic101):

2d + e = 8 d - e = 4 -----------add 3d = 12 -- divide by 3 d = 4 sub in 4 for d in either of the equations to find e d - e = 4 4 - e = 4 -e = 4 - 4 -e = 0 e = 0 solution is : d = 4 and e = 0

OpenStudy (texaschic101):

you got it right :) answer is C

OpenStudy (anonymous):

i must have did it wrong, because idk.. sory

OpenStudy (anonymous):

sweet thanks 1 more

OpenStudy (anonymous):

Use the substitution method to solve the following system of linear equations: x + y = 3 x = y – 7 A. The solution is (–2, 5). B. There is no solution. C. The solution is (4, 1). D. There are an infinite number of solutions.

OpenStudy (anonymous):

A

OpenStudy (anonymous):

r u sure -.-

OpenStudy (anonymous):

ok a i will see if it is

OpenStudy (anonymous):

ima wait to see what texaschic says

OpenStudy (anonymous):

yep... if you graph it, that is where they intersect.

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

im not sayin ur wrong but

OpenStudy (texaschic101):

lets have it :) oh..there it is... x = y - 7 sub y - 7 in for x in the other equation x + y = 3 y - 7 + y = 3 -- add 7 to both sides y + y = 3 + 7 2y = 10 y = 5 now sub 5 in for y x = y = 7 x = 5 - 7 x = -2 solution is (-2,5)

OpenStudy (anonymous):

its all good :)

OpenStudy (anonymous):

were cool B)

OpenStudy (anonymous):

good a u were right thanks u all

OpenStudy (texaschic101):

glad to help :)

OpenStudy (anonymous):

100% Assessment Summary Score: 6/6 points Status: Mastered

OpenStudy (anonymous):

sweet thanks !!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (texaschic101):

sweet :)

OpenStudy (anonymous):

before 50% 50% (3/6 points) Submitted: 12/17/13 02:13 PM Not Mastered after 100% 100% (6/6 points) Submitted: 01/17/14 09:00 AM Mastered

OpenStudy (texaschic101):

Congrads :)